OpenStudy (anonymous):

The flower garden has the shape of a right triangle. 15ft of a perennial border forms the hypotenuse of the triangle, and one leg is 3ft longer than other leg. Find the lengths if the legs. What are the lengths of the legs? ?=ft please

OpenStudy (anonymous):

Please help me someone!

OpenStudy (anonymous):

So we know the length of the hypotenuse, so call the two remaining side lengths A and B. Then, A=B+3 And, 15^2=A^2+B^2 by Pythagorean thm Solve for one of the variables, and sub into the second equation... Just ask for more help!

OpenStudy (anonymous):

I am trying to finish a few more question but I can seem to get this one worked out can I get more like the answer please by any chance

OpenStudy (anonymous):

If we substitute in the first equation into the second in terms of A, we have 15^2=(B+3)^2+B^2 So, you'll need to expand the (B+3)^2 (FOILing) and then solve for B. It'll be a quadratic so you'll need to factor, or complete the square or use the formula. Give it a try and let me know what you end up with.

OpenStudy (anonymous):

the length of the legs are 6 and 9 using the pythagorean theorem ( \[a^{2} + b^{2} = c^{2}\]) 15=c(hypotenuse) ( 3+L )^{2} + L^{2} = 15^{2} \[\sqrt{( 3+L )^{2} + L^ {2} = 15^{2}}\] 3 + L + L = 15 2L =15-3 2L = 12 L=6 for the other leg = 3+L 3 + 6 = 9 so, 6 and 9 are the lengths of the legs.:)) i hope my answer is correct..:)

OpenStudy (anonymous):

6ft and 9ft